OpenStudy (anonymous):

Simplify (8j3 – 10j2 – 7) – (6j3 – 10j2 – j + 12)

OpenStudy (anonymous):

is it 2j3 – j + 5

OpenStudy (lgbasallote):

get rid of the grouping symbol by distributing the minus so... 8j^3 - 6j^3 -10j^2 + 10j^2 + j -7 -12 is that 2j^3 - j + 5? waddyouthink? :D

OpenStudy (anonymous):

well thats what I got

hero (hero):

Looks like someone is getting the hang of this. =]

OpenStudy (anonymous):

I got 14j^3-20j^2=5 i use the kfc...

hero (hero):

You used the what?

OpenStudy (anonymous):

Keep Flip Change..its easier for me..

hero (hero):

(8j3 – 10j2 – 7) – (6j3 – 10j2 – j + 12) 8j^3 - 10j^2 - 7 - 6j^3 + 10j^2 + j - 12 8j^3 - 6j^3 - 10j^2 + 10j^2 + j - 7 - 12 2j^3 + j - 19

OpenStudy (lgbasallote):

8j^3 - 6j^3 = 2j^3 -10j^2 + 10j^2 = 0 j = j -7 - 12 = -19 so it should be....im gonna leave that to you @Winterskull1215 it seems you forgot to distribute the negative sign

OpenStudy (anonymous):

darn it..i guess i do suck at math.. :(

hero (hero):

If you put like terms together before adding or subtracting, you'd have much better results.

OpenStudy (anonymous):

i guess so.. @Hero

OpenStudy (anonymous):

I find out that I forgot a part of the question..